OpenStudy (anonymous):

find the solutions to the equation below. Check all that apply. 9x^2+9x-4=0 A.X=4 B.x=-2 C.x=-4/3 D.x=1/2 E.x=1/3 F.x=-1/3

OpenStudy (anonymous):

i got A nd C are they right

OpenStudy (anonymous):

FIrst factor out the problem... \[(3 x-1) (3 x+4) = 0 \] Then solve!

OpenStudy (anonymous):

C was correct but look at (3x -1) = 0

OpenStudy (anonymous):

it has a - sign

OpenStudy (anonymous):

Solve that equation... \[3x -1 =0\]\[3x = 1\] Whats x?

OpenStudy (anonymous):

x=1/3

OpenStudy (anonymous):

WOO-HOO!!! So they are C and E

OpenStudy (anonymous):

so its C nd E